Reports 1-1 of 1 Clear search Modify search
AdV-TCS (CO2 laser projector)
lorenzini - 12:37 Tuesday 11 June 2024 (64483) Print this report
WI CO2 laser power instability recovered

This morning, dunring the customary check tuning of CO2 laser power settings, I found the WI CO2 laser power again out of the stable range for operation, due to a drop occurred at around 6.30 UTC (see fig1 where the laser seen by the sampler is plotted). Therefore I tried to recover with the usual method of shifting the chiller setpoint, however it took some time to find a (seemingly) stable behaviour scanning temperatures from 19.04 °C up to 19.08 °C, then settling down to 19.055 °C at 9.56 UTC. The BCK chiller setpoint was changed accordingly. Then I also tuned the power settings of both NI and WI CO2 laser readings on the CH PICKOFF channel for DAS IN, DAS OUT and CH values matching the default values for O4b.

However, the readings of WI DAS IN and DAS OUT power meters are now pretty different from those before the drop. To be monitored in next days.

Images attached to this report
Comments to this report:
nardecchia, lorenzini - 23:00 Tuesday 11 June 2024 (64488) Print this report

This afternoon, between 15:20 and 15:40 UTC, we investigated the behavior of the WI CO2 laser by decreasing the chiller set point. We restored the standard behavior of the laser at 19.025 degrees (from 19.055 degrees).

Images attached to this comment
Search Help